Suppose that the demand and the total cost functions of a monopolist are P=24-3Q and C=Q2+8Q respectively, find the optimum quantity, price and profit on these levels.



1
Expert's answer
2022-02-17T14:44:49-0500

"P=MR=24-3Q"

Integrating MR yields;

"TR=24Q-\frac{3}{2}Q^2"

"TC=Q^2+8Q"

Differentiating TC yield;

"MC=2Q+8"

"MR=MC"

"2Q+8=24-3Q"

"5Q=16"

"Q=3.2 units"

Optimum quantity= 3.2units

 "P=24-3(3.2)"

"P=24-9.6"

"price=14.4 dollars"

"Profit=TR-TC"

"TR=(24(3.2)-\frac{3}{2}(3.2)^2)"

"TR=64.44 dollars"

"TC=35.84 dollars"

 "TC=3.2^2+8(3.2)=35.84"

 "Profit=64.44-35.84"


"Profit=28.6 dollars"